Wilson Co. is considering two mutually exclusive projects. Both require an initial investment of $10,000 at t = 0. Project X has an expected life of 2 years with after-tax cash inflows of $7,000 and $7,500 at the end of Years 1 and 2, respectively. In addition, Project X can be repeated at the end of Year 2 with no changes in its cash flows. Project Y has an expected life of 4 years with after-tax cash inflows of $5,600 at the end of each of the next 4 years. Each project has a WACC of 13%. What is the equivalent annual annuity of the most profitable project?

An investor bought 300 shares of Green International when it was selling for $40 a share and sold the shares one year later for $110.23. Green paid $2 per share in dividends. Calculate the investor’s rate of return.

Murray Company's bonds mature in 15 years, have a par value of $1,000, and pay an interest rate (coupon rate) of 5.5 percent annually. The market requires an interest rate of 7.2% on these bonds. What is the bond's price?
Answer #1

1. Equivalent Annual Annuity

Project X Year | Cash Flow 0 $(10,000.00)1.0000 1 7,000.00 2$ (2,500.00) 0.7831 S 3 $ 7,000.00 4$ 7,500.00 PVF @ 13% Discount

2. Investors Rate of return = (Sale Value + Dividends - Purchase Value) / Purchase value

Investors Rate of return = (33069 + 600 - 12000) / 12000

Investors Rate of return = 21669 / 12000

Investors Rate of return = 180.58%

3. Bond Price = Coupon * PVAF(7.2%, 15) + Maturity * PVF ( 7.2%, 15)

Bond Price = 55 * 8.9940 + 1000 * 0.35243

Bond Price = $847.70